economics question

It is a hot day, and Kaleem is very thirsty. Here is the value he places on a bottle of water:

Value of first bottle = Rs. 7
Value of second bottle = Rs. 5
Value of third bottle = Rs. 3
Value of fourth bottle = Rs. 1

From this information, derive Kaleem’s demand schedule. Graph his demand curve for bottled water.
If the price of a bottle of water is $4, how many bottles does Kaleem buy? How much consumer surplus does Kaleem get from his purchases? Show Kaleem’s consumersurplus in your graph.
If the price falls to $2, how does quantity demanded change? How does Kaleem’s consumer surplus change? Show these changes in your graph.
Answer #1

a)

BottlesValue of Bottles
First7
Second5
Third3
Fourth1

From the above table, the demand curve is constructed,

7 6. 5 3 Demand 2 1 Third fouth

The above represents the Kaleem Demand Curve.

b)

If Price = $4 then 2 bottles are consumed.

Cosumer Surplus = (7-4) + (5-4) = 3 + 1 = 4

P ㅋ 6 cos S S 4. 3 Demand 2 1 First seup

c)

If Price = $2 then 3 bottles are consumed.

Cosumer Surplus = (7-2) + (5-2) + (3-2) = 5 + 3 + 1 = 9

ㅋ 6 Cos S 4 3 y Demand 2 O First sewed Third Fouth

Change in Consumer surplus = 9 - 4 = 5